OpenStudy (anonymous):

x squared plus x equals 12. solve using the Quadratic Formula

OpenStudy (paki):

here... a=1, b=1, c=-12... put in quadratic formula... and solve...

OpenStudy (anonymous):

i understand that what i don't understand is how to do it

OpenStudy (paki):

put your values in this...

OpenStudy (anonymous):

then?

OpenStudy (paki):

i have putted the values... now take out calculator and solve dear...

OpenStudy (paki):

x= (-1) +- root (1^2 - 4*(1)*(-12) / 2 (1)... now solve...

OpenStudy (anonymous):

it would equal 47 right?

OpenStudy (anonymous):

or would it be 23.5?

OpenStudy (paki):

you will get 2 values,,,, x=-1+7/2 and x= -1-7/2 x= -6/2 and x= -8/2 x= -3 and x=-4

OpenStudy (anonymous):

oh..thanks paki!

OpenStudy (paki):

pleasure dear...